OpenStudy (anonymous):

what is the Slope- intercept form and the standard for for the pints (4,4) and (5,7) I got that the slope is =3

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) slope = (7 - 4) / (5 - 4) slope = 3/1 you got the slope right. Slope intercept form is y = mx + b so we use y = mx + b slope(m) = 3 use either of your points...(4,4)...x = 4 and y = 4 now we sub 4 = 3(4) + b 4 = 12 + b 4 - 12 = b - 8 = b so slope intercept form is : y = 3x - 8 standard form : Ax + By = C y = 3x - 8 --- subtract 3x from both sides -3x + y = - 8 --- multiply by -1 to make x positive 3x - y = 8 <== standard form

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

no

OpenStudy (texaschic101):

good...happy to help :)